OpenStudy (anonymous):

Could someone please explain to me step by step how to solve equations like this please Find x if 2x(squared sign) - 5x + 3 = 0

OpenStudy (anonymous):

2x^2 -2x-5x+3=0 when middle term is split

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

it is 2x^2 -2x-3x+3

OpenStudy (anonymous):

Or you can use completing the square or quadratic formula, if you have learnt those methods

OpenStudy (anonymous):

then it becomes 2x(x-1)-3(x-1)

OpenStudy (anonymous):

which will be (2x-3)(x-1)=0

OpenStudy (anonymous):

2x-3=0 and x-1=0

OpenStudy (anonymous):

x=3/2 or x=1

OpenStudy (anonymous):

@lisamc91 did you understand?

OpenStudy (anonymous):

yes I do thanks allot! :)

OpenStudy (anonymous):

you're welcome

OpenStudy (unklerhaukus):

ill show you the quadratic formula method...

OpenStudy (unklerhaukus):

Say you have some general quadratic equation \[ax^2+bx+c=0\] The solution(s) for x, can be found using the quadratic formula \[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] With the solution(s), you can factorise the original quadratic \[(x-x_1)(x-x_2)=0\]

OpenStudy (unklerhaukus):

so for your example \[2x^2 - 5x + 3 = 0\] comparing with \[ax^2+bx+c=0\] \[a=2,\qquad b=-5,\qquad c=3\]

OpenStudy (unklerhaukus):

so \[x_{1,2}=\frac{-(-5)\pm\sqrt{(-5)^2-4\times2\times3}}{2\times2}\] ______________________ ie one solution for x \[x_1=\frac{-(-5)+\sqrt{(-5)^2-4\times2\times3}}{2\times2}\] and the other \[x_2=\frac{-(-5)-\sqrt{(-5)^2-4\times2\times3}}{2\times2}\]

OpenStudy (unklerhaukus):

are you following @lisamc91 ?

OpenStudy (anonymous):

ah yes the formula really helped! Ive done a few with the formula i think im getting it now! thank you unklerhaukus! :)

OpenStudy (unklerhaukus):

( :